### v1

# X-GRM: Large Gaussian Reconstruction Model for Sparse-view X-rays to Computed Tomography

Abstract
Computed Tomography serves as an indispensable tool in clinical workflows, providing non-invasive visualization of internal anatomical structures. Existing CT reconstruction works are limited to small-capacity model architecture, inflexible volume representation, and small-scale training data. In this paper, we present X-GRM (X-ray Gaussian Reconstruction Model), a large feedforward model for reconstructing 3D CT from sparse-view 2D X-ray projections. X-GRM employs a scalable transformer-based architecture to encode an arbitrary number of sparse X-ray inputs, where tokens from different views are integrated efficiently. Then, tokens are decoded into a new volume representation, named Voxel-based Gaussian Splatting (VoxGS), which enables efficient CT volume extraction and differentiable X-ray rendering. To support the training of X-GRM, we collect ReconX-15K, a large-scale CT reconstruction dataset containing around 15,000 CT/X-ray pairs across diverse organs, including the chest, abdomen, pelvis, and tooth etc. This combination of a high-capacity model, flexible volume representation, and large-scale training data empowers our model to produce high-quality reconstructions from various testing inputs, including in-domain and out-domain X-ray projections. Project Page: https://github.com/CUHK-AIM-Group/X-GRM.
footnotetext: $\dagger$ = Corresponding Author
1 Introduction
Reconstructed from penetrative X-ray projections, Computed Tomography (CT) serves as an indispensable tool in disease diagnosis, treatment planning, and intraoperative guidance x_ray_1 ; x_ray_3 ; x_ray_4 . Traditional reconstruction methods fdk ; analytical_2 ; asd_pocs ; sart typically require several hundred X-ray projections to achieve diagnostically acceptable image quality, exposing patients to excessive radiation and also consuming substantial resources. To address these limitations, our work investigates reconstructing CT volumes from sparse-view X-ray projections.
Existing sparse-view CT reconstruction approaches can be categorized into optimization-based and regression-based methods. Optimization-based methods iteratively refine 3D volumes to match X-ray projections using neural representations zha2022naf ; shen2022nerp ; zha2024r ; cai2024radiative ; sax_nerf or diffusion models chung2023solving ; chung2023decomposed ; lee2023improving . While effective, these approaches typically require minutes to even hours for reconstructing per case, making them impractical for time-sensitive applications. Regression-based methods fbpconvnet leverage neural networks to learn useful patterns for diverse reconstruction tasks-ranging from projection extrapolation anirudh2018lose ; ghani2018deep and slice denoising wang2022dudotrans ; jin2017deep ; freeseed to direct volume regression difnet ; dif_gs ; c2rv . Despite enabling rapid inference, these approaches still face significant challenges: (1) they predominantly employ small model architectures, such as CNNs, which suffer from inherent scaling limitations, (2) they primarily represent CT volumes as 3D voxel, which does not support differentiable X-ray rendering for downstream applications, and (3) they are trained on small-scale, single-organ datasets setio2017validation ; cipriano2022deep ; mccollough2016tu , limiting their applicability to CT volumes with diverse anatomical structures.
Figure 1: Our method achieves state-of-the-art reconstruction quality while maintaining the fastest runtime. (a) Qualitative results: DIF-Gaussian dif_gs exhibits issues with over-smooth results (red boxes) and R2-Gaussian r2_gaussian has noise artifacts (orange boxes) and is time-consuming. In contrast, our method achieves better fidelity in a much shorter time. (b) Performance and runtime comparison: metrics are evaluated on the ReconX-15K test set.
To tackle these limitations, we propose a large feedforward X-ray Gaussian Reconstruction Model (X-GRM), for CT reconstruction from sparse-view X-ray projections. Our innovation is that a transformer-based architecture transformer can excel at CT reconstruction when coupled with training on large-scale datasets. Specifically, we develop a scalable X-ray Reconstruction Transformer that employs an encoder ViT to efficiently tokenize each X-ray projection in parallel, followed by a fusion ViT that integrates global information by performing self-attention across views. Meanwhile, inspired by the efficient rendering of 3D Gaussian Splatting kerbl20233d , we introduce a new volume representation, named Voxel-based Gaussian Splatting (VoxGS), where isometric 3D Gaussians are placed at voxel centers and rendering-related attributes are regressed from tokens. Such design enables efficient CT volume extraction and differentiable X-ray rendering, thereby introducing more constraints from X-ray inputs during training and also supports downstream render-related applications. Finally, to support the training of X-GRM, we introduce a large-scale CT reconstruction dataset ReconX-15K, containing approximately 15,000 volume-projection pairs with diverse anatomical structures. Unlike previous small datasets such as LUNA16 setio2017validation , which contains fewer than 1,000 cases focused on only chest regions, ReconX-15K is 16 times larger and consists of CT volumes with diverse organs, including the chest, abdomen, pelvis, and tooth etc. As shown in Fig. 1, with the proposed scalable model architecture, flexible volume representation, and large-scale data training, X-GRM drastically outperforms the SOTA methods in terms of both the reconstruction quality and inference speed.
In summary, the main contributions of our work are as follows:
  • We introduce X-GRM, a large feed-forward Transformer model that, given an arbitrary number of sparse X-ray projections, can directly predict the CT volume within one second.
  • We introduce a flexible volume representation-Voxel-based Gaussian Splatting, which allows for efficient CT volume extraction and differentiable X-ray rendering.
  • We collect a large-scale CT reconstruction dataset, ReconX-15K, containing over 16K samples of volume-projection pairs containing diverse anatomical structures.
  • We demonstrate that X-GRM drastically outperforms existing methods in terms of both the reconstruction quality and inference speed. Our codes, pre-trained models, and datasets will be available to the community.
2 Related work
2.1 Sparse-View CT reconstruction
Current approaches of sparse-view CT reconstruction fall into two fundamental categories. The first encompasses reconstruction through optimization, progressively adjusting the 3D volume to match acquired projections. Conventional approaches sart ; sauer1993local ; sidky2008image formulate this as a maximum a posteriori estimation problem, while contemporary methods incorporate neural representations zha2022naf ; shen2022nerp ; zha2024r ; cai2024radiative ; ddgs ; dgr ; sax_nerf and diffusion modeling chung2023solving ; chung2023decomposed ; lee2023improving . Though these deliver high-quality results, their substantial computational demands render them unsuitable for time-sensitive clinical applications. The second category involves direct regression methodologies using convolutional neural networks ronneberger2015u to learn patterns from extensive datasets. These systems perform reconstruction tasks including projection enhancement anirudh2018lose ; ghani2018deep , slice denoising wang2022dudotrans ; jin2017deep ; freeseed , or volume estimation difnet ; dif_gs ; c2rv . While offering significantly faster processing speeds, these approaches suffer from inherent limitations in neural architecture design, inflexible volume representation, and training on small-scale data. Our proposed X-GRM specifically addresses these fundamental challenges. Concurrent works X-LRM zhang2025x and DeepSparse lin2025deepsparse also explore large foundation models for CT reconstruction, but our work distinguishes itself through novel architecture and 3D Gaussian volume representation, enabling both efficient reconstruction and high-quality X-ray rendering from arbitrary viewpoints—a capability absent in these alternative approaches.
2.2 Feed-forward 3D reconstruction
Unlike methods based on per-sample optimization mildenhall2021nerf ; kerbl20233d , feed-forward reconstruction approaches aim to recover various 3D representations (e.g., mesh wang2018pixel2mesh ; wu2020pq , implicit fields yu2021pixelnerf , 3DGS xu2024grm ; charatan2024pixelsplat ; chen2024mvsplat , and point maps wang2024dust3r ) through a single forward pass from input image sequences. Specifically, in the domain of feed-forward object generation, LRM hong2023lrm and its variants li2023instant3d ; tochilkin2024triposr ; wang2023pf ; wei2024meshlrm ; zhang2024gs ; he2024lucidfusion ; cai2024baking have achieved significant advances in both performance and efficiency by leveraging large-scale datasets deitke2023objaverse ; deitke2023objaversexl and transformer architectures vaswani2017attention . Similarly, in feed-forward scene reconstruction, DUSt3R wang2024dust3r , pixelSplat charatan2024pixelsplat and subsequent variations mast3r ; fast3r ; mvdust3r ; flare ; cut3r ; vggt ; chen2024mvsplat ; noposplat ; wang2024freesplat ; tang2024hisplat have produced impressive results through continuous scaling up of their approaches. However, CT reconstruction remains in its infancy, lacking sufficient generalization capabilities due to the absence of large-scale reconstruction datasets and scalable models. Our work aims to fill this blank.
3 Method
3.1 Preliminaries
[…]
3D Gaussian Splatting kerbl20233d represents an approach to 3D content modeling through a collection of $N_{p}$ colored Gaussian primitives $\mathcal{G}=\{{\bm{g}}_{i}\}_{i=1}^{N_{p}}$. Each Gaussian element ${\bm{g}}_{i}=\exp\left(-\frac{1}{2}({\bm{x}}-\bm{\mu}_{i})^{\top}\bm{\Sigma}_{i}^{-1}({\bm{x}}-\bm{\mu}_{i})\right)$ is characterized by its opacity $\bm{\sigma}_{i}\in\mathbb{R}$ and color ${\bm{c}}_{i}\in\mathbb{R}^{3}$ properties, alongside positional parameters $\bm{\mu}_{i}\in\mathbb{R}^{3}$ and covariance matrix $\bm{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ that encodes scaling ${\bm{s}}_{i}\in\mathbb{R}^{3}$ and orientation quaternion ${\bm{r}}_{i}\in\mathbb{R}^{4}$ within 3D space. This comprehensive set of attributes forms ${\bm{G}}\in\mathbb{R}^{N_{p}\times 14}$, capable of encoding complex 3D scenes or objects. By aligning the formula, the rasterization process is equivalent to the CT imaging procedure, thus we can use 3DGS to represent CT volumes and render X-rays (see Sec. A.1 for more details). Our approach also modify the original 3DGS to address the distinct requirements of feed-forward CT reconstruction tasks.
3.2 Overview
Given a set of sparse-view X-ray projections =$\mathcal{I}=\{{\bm{I}}_{i}\}_{i=1}^{K},({\bm{I}}_{i}\in\mathbb{R}^{H\times W})$ and their corresponding projection matrix $\mathcal{P}=\{{\bm{P}}_{i}\}_{i=1}^{K},{\bm{P}}_{i}={\bm{K}}_{i}[{\bm{R}}_{i}|{\bm{t}}_{i}]$, calculated from via intrinsics matrix ${\bm{K}}_{i}$, rotation matrix ${\bm{R}}_{i}$, and translation vector ${\bm{t}}_{i}$, our goal is to learn a mapping $f_{\bm{\theta}}$ from X-ray projections to CT volume density field ${\bm{V}}\in\mathbb{R}^{M\times N\times L}$:
$f_{\bm{\theta}}:\{{\bm{I}}_{i},{\bm{P}}_{i}\}_{i=1}^{K}\mapsto\{{\bm{V}}(x,y,z)|x\in[1,M],y\in[1,N],z\in[1,L]\},$ (1)
where we formulate $f_{\bm{\theta}}$ as a feedforward transformer and $\bm{\theta}$ are learnable parameters optimized from a large-scale training dataset. As illustrated in Fig. 2, our method consists of a large X-ray Reconstruction Transformer to enhance the model scalability and a flexible voxel-based Gaussian Splatting that enables efficient CT volume extraction and differentiable rendering. In this section, we first describe the process of encoding X-ray inputs into compact latent tokens, integrated with camera information (Sec. 3.3). Then, we discuss the design of VoxGS and how to regress Gaussian attributes from encoded tokens (Sec. 3.4). Finally, we detailed the model training from large-scale CT collections (Sec. 3.5).
Figure 2: X-GRM is a large feed-forward transformer trained on a curated large CT reconstruction dataset. (a) X-ray Reconstruction Transformer efficiently encodes and fuses tokens from multiple X-ray projections, and (b) Voxel-based Gaussian Splatting enables both the efficient CT volume extraction and differentiable X-ray rendering.
3.3 X-ray Reconstruction Transformer
Existing CNN-based CT reconstruction methods difnet ; dif_gs ; c2rv face inherent scaling limitations that restrict their effectiveness for complex volumetric reconstruction. To overcome these constraints, we propose a large X-ray Reconstruction Transformer that significantly enhances CT reconstruction capabilities. Our approach employs an encoder ViT to efficiently tokenize X-ray projections in parallel, followed by a fusion ViT that integrates tokens across multiple views to reconstruct high-quality 3D volumes. Additionally, we incorporate camera geometry information into tokens through ModLN peebles2023scalable , enabling more precise spatial reasoning.
Encoder ViT.
We encode each X-ray projection ${\bm{I}}_{i}\in\mathcal{I}$ to a set of patch features ${\bm{H}}_{i}$, using a feature extractor $\mathcal{F}_{enc}$. This is done independently per image, yielding a sequence of image patch features ${\bm{H}}_{i}=\{h_{i,j}\}_{i,j=1}^{HW/P^{2}}$ for each image:
${\bm{H}}_{i}=\mathcal{F}_{enc}({\bm{I}}_{i}),i=1,...,K.$ (2)
Here we adopt transformer-based DINO zhang2022dino as our encoder. Before passing image features $\{{\bm{H}}_{i}\}_{i=1}^{K}$ to the decoder, we also inject Plücker ray directions into the features via adaptive layer norm peebles2023scalable :
$\tilde{{\bm{H}}_{i}}=ModLN({\bm{H}}_{i},{\bm{R}}_{i}),i=1,...,K,$ (3)
where ${\bm{R}}_{i}$ is the Plücker ray embedding of the camera angle and origin. Unlike previous works that modulate camera poses to image features using extrinsic and intrinsic matrices hong2023lrm ; liu2023zero , Plücker rays are defined by the cross product between the camera location and ray direction, offering a unique ray parameterization independent of scale, camera position and focal length.
Fusion ViT.
Most of the computation in our framework happens in the fusion ViT. We use a 16-layer transformer similar to ViT-B transformer . This fusion transformer takes the concatenated encoded image patches from all views $\{\tilde{{\bm{H}}_{i}}\}_{i=1}^{K}$ and performs all-to-all self-attention $\mathcal{F}_{fuse}$:
$\{{\bm{F}}_{1},{\bm{F}}_{2},...,{\bm{F}}_{K}\}=\mathcal{F}_{fuse}(\tilde{{\bm{H}}_{1}},\tilde{{\bm{H}}_{2}},...,\tilde{{\bm{H}}}_{K}),$ (4)
where ${\bm{F}}_{i}$ is the fused token of $i$-th view. This operation provides features of each view with full context from all other views, enables complete spatial and context reasoning.
3.4 Voxel-based Gaussian Splatting
Existing feedforward models dif_gs typically represent CT as 3D voxel, yet it has limitations of supporting differentiable real-time rendering, thus can hardly leverage the X-ray projection constraints during the training period. Benefiting from highly optimized rasterization, 3DGS kerbl20233d enables real-time differentiable rendering. However, directly representing CT with original 3DGS or modified 3DGS for CT r2_gaussian ; ddgs leads to inaccurate volume extraction and instable optimization for feedforward framework (Tab. 6(c)). Therefore, we propose a new CT representation-Voxel-based Gaussian Splatting (VoxGS) to support precise volume extraction and differentiable X-ray rendering.
Voxel-based Gaussian Splatting.
VoxGS has two different designs with the original 3DGS. Firstly, it consists of 3D Gaussians $\mathcal{G}=\{{\bm{g}}_{i}\}_{i=1}^{N_{p}}$ with fixed positions. In particular, the position $\bm{\mu}_{i}$ of each Gaussian ${\bm{g}}_{i}$ is fixed at the centroid of each voxel grid ${\bm{V}}(x,y,z)$:
$\bm{\mu}_{i}=(x,y,z),x\in[1,M],y\in[1,N],z\in[1,L].$ (5)
This design offers the advantage that when extracting CT volumes from 3D Gaussians $\mathcal{G}$, complex trilinear interpolation strategies are not required for identifying the voxel value, instead we can directly obtain it from the Gaussians opacities, which improves extraction speed and enables more stable optimization. Secondly, we remove the SH coefficients ${\bm{c}}_{i}$ related to color rendering from the Gaussian attributes, retaining only the opacity $\alpha_{i}$, the scale vector ${\bm{s}}_{i}$, and the rotation vector ${\bm{r}}_{i}$. This design is similar to r2_gaussian , where the kernel formulation removes view-dependent color because X-ray attenuation depends only on isotropic density. In summary, we can represent a CT volume as a set of voxel-based Gaussians: $\mathcal{G}=\{{\bm{g}}_{i}|{\bm{g}}_{i}=\{\bm{\mu}_{i},\alpha_{i},{\bm{s}}_{i},{\bm{r}}_{i}\},\bm{\mu}_{i}\text{ is fixed}\}_{i=1}^{N_{p}}$, from which X-ray projections can be efficiently rendered in a differentiable manner.
Decoding VoxGS attributes.
As Gaussian positions $\bm{\mu}_{i}$ is fixed, we only need to decode Gaussians attributes $\alpha_{i}$ and $\bm{\Sigma}_{i}$ from available view features $\{{\bm{F}}_{1},{\bm{F}}_{2},...,{\bm{F}}_{K}\}\subset\mathbb{R}^{H\times W\times C}$, where $K$ is the number of X-ray projections. Specifically, for a 3D Gaussian ${\bm{g}}_{i}$ with fixed position $\bm{\mu}_{i}$, we query view-specific features $\tilde{{\bm{F}}}_{k}$ from ${\bm{F}}_{k}$:
${\bm{F}}_{k}^{i}=Interp({\bm{F}}_{k},{\bm{P}}_{k}(\bm{\mu}_{i}))\in\mathbb{R}^{C},\text{ for }k=1,2,...,K,$ (6)
where ${\bm{P}}_{k}:\mathbb{R}^{3}\rightarrow{\mathbb{R}^{2}}$ is the projection matrix of $k$-th view, and $Interp(\cdot)$ denotes bilinear interpolation. Then, $K$ queried features are aggregated by a MaxPooling layer to obtain ${\bm{F}}^{i}_{max}=MaxPool({\bm{F}}_{1}^{i},{\bm{F}}_{2}^{i},...,{\bm{F}}_{K}^{i})\in\mathbb{R}^{C}$. Finally, ${\bm{F}}^{i}_{max}$ is passed to several MLP layers to predict Gaussian parameters:
$[\alpha_{i},{\bm{s}}_{i},{\bm{r}}_{i}]=MLPs({\bm{F}}_{max}^{i})\in\mathbb{R}^{1+3+4}.$ (7)
Volume extraction and X-ray rendering.
To extract the CT density volume ${\bm{V}}\in\mathbb{R}^{M\times N\times L}$, we can directly index the corresponding fixed Gaussian opacities $\alpha_{i}$:
${\bm{V}}(x,y,z)=\alpha_{i},\text{ where }\bm{\mu_{i}}=(x,y,z)^{T}.$ (8)
Benefiting from the fixed position design of VoxGS, extracting the density field can be achieved through fast indexing operations, without need of extra computation. While for X-ray rendering, we employ the rasterizer implemented in r2_gaussian to render projections in a differentiable manner, which enables more constraints from inputs X-rays and also support various downstream applications, like novel view synthesis and potential CT/X-ray registration.
3.5 Training objectives and protocols
Training objectives.
We train the network with two objectives: 1) to impose constraints on the predicted CT volume ${\bm{V}}$ such that it aligns with the ground truth ${\bm{V}}_{gt}$ and 2) to minimize the difference between the actual X-ray projections ${\bm{I}}_{gt}$ and rendered ones ${\bm{I}}$. Accordingly, we adopt two categories of losses: volume constraints and rendering constraints: $\mathcal{L}=\mathcal{L}_{volume}+\mathcal{L}_{render}$. Specifically, for volume reconstruction constraint $\mathcal{L}_{volume}$, we adopt MSE to compute point-wise estimation error:
$\mathcal{L}_{volume}({\bm{V}},{\bm{V}}_{gt})=\sum_{x,y,z}({\bm{V}}(x,y,z)-{\bm{V}}_{gt}(x,y,z))^{2}.$ (9)
While for rendering constraint $\mathcal{L}_{render}$, we use a weight combination of photometric L1 loss $\mathcal{L}_{1}$ and D-SSIM loss $\mathcal{L}_{ssim}$ ssim :
$\mathcal{L}_{render}({\bm{I}},{\bm{I}}_{gt})=\lambda_{L1}\mathcal{L}_{1}({\bm{I}},{\bm{I}}_{gt})+\lambda_{ssim}\mathcal{L}_{ssim}({\bm{I}},{\bm{I}}_{gt}).$ (10)
Training protocols.
During the training period, we observe rendering with the original volume resolution $M\times N\times K$ ($256^{3}$ in this work), requires exhaustive GPU memory to preserve gradients. Therefore, we instead randomly sample sub-volumes $\tilde{{\bm{V}}}\in\mathbb{R}^{M/4\times N/4\times K/4}$ from the original volume ${\bm{V}}$, and modify the training objectives of Eq. 9 accordingly. While during the inference, we directly use the full volume ${\bm{V}}$, as no gradients are required to store.
Table 1: The statistics of ReconX-15K.
Dataset Body Regions # Volumes
AbdomenAtlas v1.0 li2024abdomenatlas Abdomen, Chest, Pelvis 5,171
RSNA2023 hermans2024rsna Abdomen, Pelvis 4,711
LUNA16 setio2017validation Chest 833
AMOS ji2022amos Abdomen 1,851
MELA mela_dataset Chest 1,100
RibFrac ribfracclinical2020 ; ribfracchallenge2024 Abdomen, Chest 660
ToothFairy2 toothfairy Tooth 223
STSTooth ststooth Tooth 423
ReconX-15K (Ours) All above 14,972
4 Experiment
4.1 Experimental setup
Datasets.
Existing CT reconstruction datasets cipriano2022deep ; mccollough2016tu ; setio2017validation exhibit two limitations: restricted data size (<1,000 cases) and single-organ specificity, thereby compromising the potential for developing generalizable and robust models. Addressing this research gap, we establish ReconX-15K, a large and diverse dataset containing 14,972 CT volumes and their corresponding X-ray projections curated from 8 public datasets (detailed in Tab. 1). It spans the anatomical regions most relevant to clinical applications, covering chest, abdomen, pelvis, and dental structures. To facilitate model development and evaluation, we partition ReconX-15K into three subsets designated as 13,612, 680, and 680, serving as training, validation, and testing sets respectively. Detailed split of each subset is describe in Sec. A.2.
To standardize CT volumes, we resample anatomical regions to consistent dimensions: chest, abdomen, and pelvis volumes to $50^{3}\text{cm}^{3}$ with $256^{3}$ voxels; dental volumes to $40^{3}\text{cm}^{3}$ with $256^{3}$ voxels. Radiodensity values are normalized from Hounsfield units ([-1000, 1000] for chest, abdomen, pelvis; [-1000, 3000] for dental structures) to the range [0,1], ensuring coverage of primary organs of interest. As most public datasets provide only CT volumes, we generated multi-view projections using the TIGRE biguri2016tigre toolbox. For each volume, we rendered 50 X-ray projections at $256^{2}$ resolution, evenly distributed across the full $[0^{\circ},360^{\circ}]$ range. To enhance realism, we incorporated Gaussian and Poisson noise to simulate Compton scattering effects.
Implementation details.
The proposed X-GRM is implemented with PyTorch pytorch and trained on 4 NVIDIA A100 (40G) GPUs. The model is optimized with AdamW loshchilov2017adamw for 100 epochs, with a batch size of 8 and $\beta_{1}=0.9,\beta_{2}=0.95$. The initial learning rate is set to $1\times 10^{-4}$ and gradually decreases to $1\times 10^{-6}$ following the cosine annealing scheduler loshchilov2016sgdr . For the model implementation, the encoder ViT $\mathcal{F}_{enc}$ uses a ViT-B/16 transformer architecture, initialized from DINO zhang2022dino pre-trained weights. The fusion ViT $\mathcal{F}_{fuse}$ is a ViT-B/16 model initialized from scratch, which has $16\times 16$ patch size, 12 layers, 12 heads, embedding dimension 768, and MLP ratio 4.0. We leverage bFloat16 precision and gradient checkpointing to improve GPU memory and computational efficiency.
During model training, we employ a variable-view learning strategy with the number of views randomly selected from $\{6,8,10\}$, and views are evenly sampled from the 50 available projections. For quantitative evaluation, we utilize both PSNR (calculated directly in 3D volumetric space) and SSIM ssim (averaged across 2D slice comparisons) to comprehensively assess reconstruction quality through signal fidelity and structural similarities.
Table 2: Quantitative comparison with traditional and feedforward methods. We evaluate all methods using the full ReconX-15K test set (2563 resolution) on one A100 GPU. All feedforward models are trained on the same training set. Bold: best results; underlined: second-best results.
       Method     Time (s)$\downarrow$            6-View            8-View            10-View
[…]
FDK fdk 0.297 14.65 0.297 15.50 0.295 16.23 0.314
ASD-POCS asd_pocs 5.704 21.47 0.498 22.32 0.560 22.62 0.653
SART sart 21.091 21.89 0.634 22.52 0.642 23.66 0.685
FBPConvNet fbpconvnet 0.131 22.63 0.691 23.05 0.695 23.48 0.718
FreeSeed freeseed 0.508 24.67 0.781 24.89 0.789 25.43 0.790
DIF-Net difnet 1.561 22.27 0.684 22.72 0.688 23.27 0.714
DIF-Gaussian dif_gs 2.120 24.83 0.787 25.03 0.793 25.58 0.794
X-GRM (Ours) 0.928 28.39 0.873 28.86 0.879 29.21 0.886
Figure 3: Qualitative comparison with traditional and feedforward methods. Results shown are from the test set reconstructions with 10-view inputs.
4.2 Assessing CT reconstruction
We evaluate our X-GRM against existing methods with two settings. Setting 1: evaluating on the full test set (680 samples) of ReconX-15K. This setting includes traditional and feedforward methods, as they can efficiently handle massive X-ray inputs. To ensure fair comparison, we re-train the compared feedforward models on the same training set. Setting 2: evaluating on a sampled set (40 samples). This setting includes self-supervised approaches, as they typically require substantial optimization time. To cover diverse organs, we evenly sample five volumes from the test set of eight datasets.
Setting 1. We first compare with traditional methods (FDK fdk , SART sart , ASD-POCS asd_pocs ) and feedforward methods (FBPConvNet fbpconvnet , FreeSeed freeseed , DIF-Net difnet , DIF-Gaussian dif_gs ). As shown in Tab. 2, X-GRM outperforms the compared models in both reconstruction quality and inference speed. When compared to FreeSeed, the SOTA 2D feedforward model, X-GRM achieves substantial improvements of 3.71dB, 3.97dB, and 3.78dB in PSNR for 6-view, 8-view, and 10-view inputs, respectively. Similarly, X-GRM outperforms DIF-Gaussian, the leading 3D feedforward method, by 3.56dB, 3.83dB, and 3.63dB while delivering 2×2$\times$2 × inference speed. Some qualitative results are presented in Fig. 4 (more results in Sec. A.6). Traditional methods like SART introduce prominent streak artifacts in sparse-view reconstructions. Meanwhile, feedforward approaches like FreeSeed and DIF-Gaussian suffer from significant blurring, particularly in bone structures and regions with low tissue contrast. In contrast, our proposed X-GRM method can preserve fine anatomical details while maintaining structural integrity.
Setting 2. We then compare with self-supervised methods (NAF zha2022naf , SAX-NeRF sax_nerf , R2-Gaussian zha2022naf ). As shown in Tab. 8, X-GRM has clear advantages over existing self-supervised methods: it achieves the best reconstruction performance while requiring the least computational time. Compared to the second-fastest self-supervised method NAF, X-GRM is 200$\times$ faster, requiring only 0.9 second to reconstruct one CT volume of resolution 2563. While compared to the state-of-the-art method SAX-NeRF and R2-Gaussian, X-GRM delivers 4.78/4.34/3.85dB and 4.81/4.01/3.14dB higher PSNR when processing 6/8/10 inputs views. As evident from the reconstructed slices in Fig. 4 (more results in Sec. A.6), NAF exhibit noise artifacts, SAX-NeRF produces less noise, but many anatomical structures are excessively smoothed, R2-Gaussian better recovers anatomical structures, but exhibits many discontinuous regions, likely due to floating artifacts from 3DGS. In contrast, the reconstructed slices from X-GRM exhibit clearer anatomical structures, more accurate bone density, and fewer artifacts throughout the volume slices.
Table 3: Quantitative comparison with self-supervised methods. We evaluate all methods using 40 CT volumes ($256^{3}$ resolution) sampled from the full test set on one RTX 4090Ti GPU. Bold and underlined values indicate best and second-best results.
[…]
NAF zha2022naf 3.0m 20.69 0.530 21.93 0.563 22.75 0.581
SAX-NeRF sax_nerf 48.5m 22.93 0.686 23.82 0.702 24.62 0.730
R2-Gaussian r2_gaussian 8.5m 22.90 0.662 24.09 0.689 25.33 0.727
X-GRM (Ours) 0.9s 27.71 0.838 28.16 0.846 28.47 0.854
Figure 4: Qualitative comparison with self-supervised models. Results shown are from the test set reconstructions with 10-view inputs.
Table 4: Cross-dataset reconstruction evaluation.
Method Time FUMPE masoudi2018 PENGWIN liu2023pelvic
PSNR SSIM PSNR SSIM
FreeSeed freeseed 0.5s 22.69 0.760 25.09 0.824
DIF-Net difnet 1.5s 20.49 0.640 23.72 0.724
DIF-Gaussian dif_gs 2.0s 23.50 0.768 25.46 0.827
SAX-NeRF sax_nerf 48.5m 23.58 0.722 24.72 0.745
R2-Gaussian r2_gaussian 8.6m 23.09 0.679 26.19 0.782
X-GRM (Ours) 1.0s 24.81 0.811 26.04 0.856
4.3 Assessing cross-dataset generalization
Due to the use of scalable architecture and large-scale training, X-GRM is inherently superior in generalizing to out-of-distribution inputs. To demonstrate this advantage, we conduct two cross-dataset evaluations. In particular, we use the models trained on ReconX-15K, and directly test them on unseen PENGWIN liu2023pelvic (pelvis) and FUMPE masoudi2018 (chest) datasets. As evident in Tab. 4, our X-GRM significantly outperforms other feedforward models, including FreeSeed, DIF-Net, and DIF-Gaussian. Compared to self-supervised models that uses per-sample optimization, our method also shows comparable or better performance, while delivering much faster inference speed (around 500× faster compared to R2-Gaussian).
Table 5: Novel view synthesis results. Time: rendering time of one X-ray projection.
Method PSNR $\uparrow$ SSIM $\uparrow$ Time(s) $\downarrow$
NAF 28.29 0.505 0.448
SAX-NeRF 31.85 0.659 0.916
R2-Gaussian 38.26 0.955 0.003
X-GRM (Ours) 49.44 0.993 0.020
Figure 5: Qualitative results of novel views.
4.4 Assessing novel X-ray synthesis
Novel X-ray synthesis enables clinicians to examine anatomical structures, enhancing diagnostic capabilities and also facilitating dataset augmentation. We evaluate X-GRM against established self-supervised methods that support X-ray synthesis using NeRF or 3DGS representations. Our evaluation is conducted on the test set comprising 30 distinct CT samples. For each sample, we use 10 projection views as input, and evaluate using the rest 40 unseen novel views. As reported in Tab.5, X-GRM delivers 12.55dB higher PSNR over R2-Gaussian while maintaining fast rendering speeds of 0.02s for each projection. Qualitative analysis in Fig.5 shows that SAX-NeRF and R2-Gaussian generate substantial noise artifacts at tissue-air interfaces (highlighted in red boxes), whereas X-GRM consistently preserves clear, well-defined anatomical boundaries.
4.5 Ablation study
Camera pose integration. We assess the effect of camera pose integration with two variants, ‘w/o pose’ that removes the pose integration, and ‘dense add’ that directly add the ray embeddings to the encoded tokens. As shown in Tab. 6(c) (a), removing camera pose leads to 0.28db decline in PSNR, indicating pose information is necessary for reasoning CT geometry. Besides, ‘dense add’ delivers slightly worse results than ModLN, so we use ModLN as the final design.
Volume representation. We evaluate two alternative approaches: ‘w/o VoxGS’ that uses the voxel grid as the CT representation, and ‘VoxGS w/ shift’ that introduces a shift attribute to predict Gaussian positions, where CT volumes are extracted by querying from nearby 3D Gaussian distributions. As reported in Tab. 6(c) (b), the ‘w/o VoxGS’ variant delivers 0.55dB lower PSNR, suggesting that the rendering constraints provided by the Gaussian representation benefit model learning. More importantly, the ‘VoxGS w/ shift’ variant delivers a performance drop of 2.37dB in PSNR. This likely stems from the inherent difficulty in optimizing precise Gaussian positions, and inaccuracies introduced during volume extraction.
Cross-view aggregation. We conduct ablation studies on cross-view aggregation strategies, evaluating ‘w/o attention’ which eliminates the fusion ViT component, and ‘cross-attention’ which substitutes self-attention in the fusion ViT with cross-attention mechanisms. As detailed in Tab. 6(c) (c), removing attention results in a performance decline of 0.52dB in PSNR, confirming that effective cross-view information aggregation is essential for geometric reasoning. Meanwhile, the ‘cross-attention’ variant achieves comparable results to our self-attention implementation. Based on these findings, we adopt the self-attention approach for its implementation simplicity.
Table 6: Ablation study of component designs on the ReconX-16K dataset with 10-view inputs.
(a) Camera pose integration.
[…]
w/o pose 28.93 0.862
dense add 29.15 0.882
Ours (ModLN) 29.21 0.886
(b) Volume representation.
[…]
w/o VoxGS 28.66 0.858
VoxGS w shift 26.84 0.837
Ours (VoxGS) 29.21 0.886
(c) Cross-view aggregation.
[…]
w/o attention 28.79 0.864
cross-attention 29.18 0.882
Ours (Self-attn) 29.21 0.886
5 Conclusion
This paper presents X-GRM, a large feedforward model for reconstructing 3D CT volumes from sparse-view X-rays using a scalable transformer architecture and Voxel-based Gaussian Splatting representation. Trained on the collected large ReconX-15K dataset, X-GRM demonstrates superior reconstruction quality with varied number of X-ray inputs. By addressing previous limitations in model capacity, volume representation, and training data scale, our work represents an advancement in non-invasive sparse-view CT imaging for clinical applications.
Appendix A Technical Appendices and Supplementary Material
A.1 Represent CT volumes as 3D Gaussians
In 3D Gaussian Splatting (3DGS) methods [29] for natural scene rendering, the attenuation of light as it passes through a medium is described using an exponential transmittance model $T(t)$. This model is physically based on the Beer-Lambert law [60], representing the light propagation process from one point ${\bm{r}}(t_{0})$ to another point ${\bm{r}}(t)$ in space:
[…]
Here, ${\bm{r}}(t)={\bm{o}}+t{\bm{d}}$ represents a 3D point along a ray with origin ${\bm{o}}$ and direction ${\bm{d}}$, while $\sigma({\bm{p}})$ denotes the density at point ${\bm{p}}$. As established in Sec. 3.1, X-ray attenuation follows an identical model. Consequently, 3DGS rendering can be directly leveraged for X-ray projection rasterization under a simplified imaging model that accounts solely for isotropic absorption (per Beer-Lambert law), with optional exclusion of color attributes.
A.2 Dataset details
As stated in the main text, we collect 8 public datasets covering organs and structures including chest, abdomen, pelvis, and teeth. We apply data standardization to adjust the data through three steps: resampling spacing, adjusting resolution to 2563, clipping HU values, and normalizing to $[0,1]$. Then, we split each dataset into training/validation/test splits with ratio 20/1/1. Meanwhile, as most datasets solely contain CT volumes, we utilize the TIGRE toolbox [3] to render X-ray projections. The specific parameter settings for each dataset are shown in the Tab. 7.
Table 7: The parameters used to pre-process each dataset of ReconX-15K.
Dataset # Volumes Train/Val/Test Spacing (mm) Clipping range
AbdomenAtlas v1.0 [33] 5,171 4701/235/235 [1.0,1.0,1.0] [-1000, 1000]
RSNA2023 [21] 4,711 4283/214/214 [1.0,1.0,1.0] [-1000, 1000]
LUNA16 [52] 833 757/38/38 [1.0,1.0,1.0] [-1000, 1000]
AMOS [25] 1,851 1683/84/84 [1.0,1.0,1.0] [-1000, 1000]
MELA [47] 1,100 1000/50/50 [1.0,1.0,1.0] [-1000, 1000]
RibFrac [28, 73] 660 600/30/30 [1.0,1.0,1.0] [-1000, 1000]
ToothFairy2 [4] 223 203/10/10 [0.3,0.3,0.3] [-1000, 2000]
STSTooth [67] 423 385/19/19 [0.3,0.3,0.3] [-1000, 2000]
A.3 Additional experiments of novel view synthesis
Due to space constraints, Tab. 5 only presents results for novel view synthesis with 10 input views. Here, we further compare performance with 6 and 8 input views, as shown in the table. Our method consistently outperforms existing self-supervised approaches by a significant margin. In particular, compared to the current best method R2-Gaussian, our approach delivers 14.92/13.01/11.18 dB improvement in performance with 6/8/10 input views, respectively. Meanwhile, our method achieves a rendering speed of 0.02s for each projection, meeting the requirements for most scenarios. We also provide more qualitative comparisons of these methods in Fig. 6, from which we can observe that NAF fails to render correct structures, SAX-NeRF and R2-Gaussian are prone to deliver artifacts. In contrast, our X-GRM can deliver obvious better rendering quality with more accurate bone density and clearer anatomical structures.
Table 8: Additional experiments of novel view synthesis on the sampled test set (40 samples).
       Method     Render Time (s)$\downarrow$            6-View            8-View            10-View
[…]
NAF [80] 0.448 23.11 0.404 26.07 0.463 28.29 0.505
SAX-NeRF [6] 0.916 27.72 0.562 29.09 0.603 31.85 0.659
R2-Gaussian [79] 0.003 34.23 0.912 36.33 0.930 38.26 0.955
X-GRM (Ours) 0.020 49.15 0.991 49.34 0.992 49.44 0.993
Figure 6: Additional qualitative comparison of novel view synthesis.
Figure 7: Qualitative comparison of reconstructed slices with traditional and feedforward models.
Figure 8: Qualitative comparison of reconstructed slices with self-supervised models.
A.4 Limitations
X-GRM exhibits several limitations that remain future investigation. The model incurs increased memory consumption when utilizing excessive numbers of 3D Gaussians, and demonstrates suboptimal performance when reconstructing from extremely sparse inputs, particularly single X-ray projections. These challenges in future work through the implementation of efficient Gaussian pruning strategies and the integration of generative modeling approaches, which we leave as future works.
A.5 Broader impacts
Impacts on real-world applications. X-GRM delivers superior reconstruction quality with faster inference speed, enabling efficient clinical deployment. Its accuracy and performance support time-sensitive diagnostic workflows while reducing patient radiation exposure. The VoxGS representation can potentially enable many downstream applications like CT/X-ray registration (as done in DDGS-CT [18]), extending the model’s utility beyond reconstruction to integrated clinical pipelines.
Impacts on research community. The ReconX-15K dataset establishes a comprehensive benchmark for CT reconstruction methods, enabling standardized evaluation and fair comparison. This work demonstrates the transformative potential of large-scale data and high-capacity models in medical imaging, encouraging similar approaches in related domains. By open-sourcing code, models, and data, X-GRM contributes essential resources for advancing medical image reconstruction research.
A.6 More visualization
We provide more qualitative comparisons with traditional, feedforward, and self-supervised methods in Fig. 7 and Fig. 8. Our X-GRM consistently delivers better reconstruction quality with clearer anatomical structures and less noise artifacts.